OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4 + 16x2

hartnn (hartnn):

write x^4 as x^2*x^2....then take out x^2 common...can u do this?

OpenStudy (anonymous):

take x2 common

OpenStudy (anonymous):

f(x) = x^2 (x^2 + 16)

OpenStudy (anonymous):

there you go! (:

OpenStudy (anonymous):

I got f(x) = x2 (4x + i)(4x - i)

OpenStudy (anonymous):

hmm, I didn't know that you were asking to factor so far. Are there answer choices?

hartnn (hartnn):

its should be f(x) = x2 (x + 4i)(x - 4i) check again...

OpenStudy (anonymous):

How did you get (x + 4i)(x- 4i) ? @hartnn

hartnn (hartnn):

do u know what 'i' stands for??

OpenStudy (anonymous):

Imaginary number. A negative radical. i = -1

hartnn (hartnn):

its an imaginary number,but its square root of -1

hartnn (hartnn):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

I know...

OpenStudy (anonymous):

So how did you find the sum of squared roots?

hartnn (hartnn):

ok,so \[\sqrt{a^2+b^2}=(a+i*b)(a-i*b)\] i have used this...now it would be clear to u...

OpenStudy (anonymous):

Thanks (: I gave you a medal! @hartnn

hartnn (hartnn):

@gps she asked whether its f(x) = x2 (4x + i)(4x - i) which means she knew it... @Lol9999 thanks